# Advancing Reference-free Evaluation of Video Captions with Factual Analysis

### v1

@@ Abstract
Video captions offer concise snapshots of actors, objects, and actions within a video, serving as valuable assets for applications such as question answering and event localization. However, acquiring human annotations for video captions is costly or even impractical, especially when dealing with diverse video domains. Existing models trained on supervised datasets face challenges in evaluating performance across different domains due to the reliance on reference-based evaluation protocols, which necessitate ground truth captions. This assumption is unrealistic for evaluating videos in the wild. To address these limitations, we propose a reference-free evaluation framework that does not require ground truth captions, focusing on factual grounding to ensure accurate assessment of caption quality. We introduce VC-Inspector, a novel caption quality evaluator that is both reference-free and factually grounded. Utilizing large language models, we generate pseudo captions of varying quality based on supervised data, which are subsequently used to train a multimodal model (i.e., Qwen2.5-VL) as the evaluator. Our approach demonstrates superior alignment with human judgments on the VATEX-Eval dataset, outperforming existing methods. The performance also generalizes to image caption datasets, Flickr8K-Expert and Flickr8K-CF, when viewing images as 1-frame videos. Overall, VC-Inspector offers a scalable and generalizable solution for evaluating the factual accuracy of video captions, paving the way for more effective and objective assessment methodologies in diverse video domains.
@@ 1 Introduction
Figure 1: Existing reference-free metrics like EMScore [31] often fail to detect factual inaccuracies and lack a consistent scoring scale. VC-Inspector addresses these limitations by providing factually grounded, interpretable evaluations with explanations.
Video captioning plays a pivotal role in bridging visual content with natural language, offering concise descriptions of salient entities, actions, and interactions within a video. These captions are instrumental in enabling a wide range of downstream applications, including video-based question answering [27], event localization [28, 12], and content retrieval [8]. To ensure the quality and reliability of captions, robust evaluation metrics are vital for advancing the field of video-language research. However, current captioning models, predominantly trained on supervised datasets [14, 41], rely heavily on reference-based evaluation protocols [26, 3, 19, 33, 39] that compare generated captions against human-written ground truth. While effective in controlled settings, these metrics face significant limitations in real-world scenarios. First, generating such references is prohibitively labor-intensive and costly, often becoming unavailable or infeasible to obtain when scaling across diverse video domains with varying visual and contextual characteristics. Second, these metrics demonstrate a limited ability to understand the semantic meanings of the captions and often fail to recognize valid alternatives.
The challenges of assessing open-domain or in-the-wild videos necessitate the development of evaluation protocols that do not rely on reference captions. However, this remains underexplored within the video research community. Current reference-free metrics [31, 30] typically measure the visual-language semantic alignment using pretrained multimodal embeddings [29], with the comprehension of caption content restricted by the context length of the underlying text encoder. In addition, these metrics frequently overlook factual inaccuracies in the candidate captions, as shown in Figure 1. Furthermore, they are primarily image-based solutions, making them suboptimal for video content.
To address these limitations, we propose a novel reference-free evaluation framework, VC-Inspector, that eliminates the dependency on human-annotated captions. Our approach emphasizes factual grounding, ensuring that the evaluation of caption quality is based on the alignment between the caption and the actual video content, rather than comparison to a predefined reference. We build VC-Inspector atop a lightweight large multimodal model (LMM), equipping the model with the ability to judge the quality of video captions based on factual accuracy, through instruction tuning. However, while supervised video caption datasets are already scarce, captions of imperfect quality are even harder to obtain. We tackle this challenge by introducing a novel data generation pipeline powered by a large language model (LLM), creating pseudo captions of varying quality by altering the factual elements (i.e., objects and actions) in the ground truth captions from ActivityNet [14]. This assembles ActivityNet-FG-It, comprising 44K data points for instruction tuning. Figure 2 illustrates the data generation and training pipeline of VC-Inspector. Unlike previous metrics that offer only a single quality score, our approach also requires the model to explain the rationale behind its scoring, making the evaluator more interpretable.
@@
We start with evaluating VC-Inspector on the ActivityNet-FG-Eval and YouCook2-FG-Eval datasets, where the labels are synthetically generated with the same data generation pipeline as the instruction tuning dataset. Results demonstrate that the quality estimates are consistent across datasets, and VC-Inspector is versatile for different visual content and caption lengths. We further evaluate the correlation between the quality scores given by VC-Inspector and human judgments using the VATEX-Eval dataset [31], a standard benchmark for video captioning metric evaluation. The proposed evaluator stands out compared to the previous reference-free metrics, even surpassing most of the reference-based metrics. Finally, we extend the evaluation to image caption datasets, Flickr8K-Expert and Flickr8K-CF [10], showing the generalization of the proposed metric.
Overall, this work makes the following contributions:
  • We present one training dataset and two evaluation datasets consisting of ground truth captions, synthetic candidate captions, and pseudo quality scores, generated using a novel data generation pipeline.
  • We propose VC-Inspector, an open-source reference-free evaluator featuring factual grounding, with a text context length of 32K tokens and support for long videos.
  • Our evaluator outperforms existing reference-free metrics for video captions, achieving higher correlations with human evaluations and generalizing across visual domains.
@@ 2 Related Works
Traditional text-based evaluation can be extended to video captioning by discarding the video and treating the reference caption as ground truth. Alternatively, image-caption metrics can be adapted by sampling images from the video and aggregating image-caption scores. We review these metrics and their limitations below:
@@
Several rule-based metrics have been developed to evaluate generated texts, including widely used methods such as BLEU [26], METEOR [3], ROUGE [19], and CIDEr [33]. While these metrics perform reasonably well in controlled settings, they primarily focus on syntactic structure and fail to capture the semantic meaning of the generated text. To address this limitation, the SPICE series [1, 18] attempts to embed semantic understanding by parsing both reference and candidate captions into objects, attributes, and their relationships. However, these metrics still struggle to account for semantically similar or identical words expressed differently. This persistent challenge has motivated the development of embedding-based evaluation methods, such as BERTScore [40] and its extended version [37], which consider the intrinsic variance between multiple ground truth captions. More recently, CLAIR [5] explores an LLM-as-a-Judge approach for image caption evaluation, demonstrating stronger correlation with human judgments than the above metrics. However, these metrics all rely exclusively on reference captions for evaluation and do not incorporate the visual input that the captions are meant to describe.
@@
To address the limitation of text-only metrics, various approaches have been explored that cross-reference between visual input and captions. VIFIDEL [24] computes semantic similarity by matching object names extracted from the image with those mentioned in the candidate caption. However, objects are discrete representations limited to fixed categories and cannot capture the motion dynamics that are particularly informative for videos. Visual-language models (VLMs), by contrast, provide continuous representations that offer richer semantic alignment between visual and textual modalities. ViLBERTScore [16] extends BERTScore by leveraging the pretrained ViLBERT [23] models, but still requires comparison to reference captions. UMIC [15] and CLIP-based metrics [9, 13, 7, 30, 34, 38, 17, 31, 21] relax this constraint by employing contrastive learning to directly model the semantic alignment between the images and captions. Although promising for short captions, these embedding-based methods are constrained by the context length limitations of their text encoders. Recently, the strong contextual reasoning and instruction following capability of LLMs have been leveraged for evaluation across diverse tasks [22]. Building on this trend, Maeda et al. [25] employ an LLM to compare the candidate caption to the visual context extracted with a VLM. However, these methods only consider static images and do not model the temporal dynamics of video, resulting in suboptimal performance on video caption evaluation.
@@
In summary, text-based methods require costly, high-quality reference captions, while image-based extensions are suboptimal for video content. These challenges highlight the growing demand for reference-free video caption evaluation tailored for video inputs. However, it remains an underexplored topic in the community, compared to text-based and image-caption evaluation. EMScore [31] was an early attempt that supports evaluating video captions without a reference. Although it considers both frame-level and video-level embeddings, these embeddings are still obtained from an image-based encoder [29], and the text encoder is limited by a short context length. PAC-S [30] and FactVC [21] augment EMScore with positive and negative data synthesis, respectively. While they share some similarity to this work, they consider only a single level of corruption, with binary (positive/negative) differentiation, whereas our method incorporates captions with varying degrees of quality, enabling a more nuanced evaluation. In addition, these metrics produce a single scalar score without offering explanations of their judgments, posing challenges for interpreting the quality evaluation. More recently, G-VEval [32] extended the G-Eval [22] approach to video by stitching together three frames from each video. However, it is unclear whether this image-based prompting generalizes to longer or more dynamic videos. Moreover, the dependence on proprietary models such as GPT-4o limits its scalability and practicality for widespread use. In contrast, VC-Inspector finetunes a lightweight, open-source LMM to produce both evaluation scores and reasoning, offering a more interpretable and scalable solution.
Ground Truth Candidate Caption Rouge-L SPICE Soft-SPICE
The man is feeding a cat on the sofa in the living room The man is feeding a lion on the sofa in the living room 92.31 75.00 78.57
The girl is dancing in the room The girl is sleeping in the room 85.71 66.67 88.19
Table 1: Text-based metrics rely on ground truth for evaluation and often fail to detect factual errors in objects/actions (bolded).
3 Video Caption Quality Estimation
In this section, we provide a concise overview of existing metrics used for evaluating video captions, identify their shortcomings, and outline our objectives.
@@ 3.1 Overview
3.2 Toward Factually Grounded Reference-free Evaluation
In this paper, we aim to develop a factually grounded and reference-free video caption quality evaluator that can generate scores on an absolute scale, and is versatile for various domains and text lengths. We argue that an evaluation protocol is only reliable if it is factually grounded, meaning that the quality assessment must accurately reflect the correctness of object entities and actions in the caption with respect to the video. For example, the evaluator should assign a lower score to a candidate caption where an object is missing or an action is wrong in the video than a correct one, and an even worse score for a caption that has more errors. However, it remains unclear whether existing metrics can effectively capture these subtle changes in factual elements. To investigate this, we probe existing metrics with candidate captions that contain incorrect objects or/and actions. Table 1 and Figure 1 illustrate such cases for reference-based and reference-free metrics, respectively. These results reveal that current metrics are insensitive to these factual errors. Even when captions are semantically distinct (e.g., “cat" vs. “lion" and “dancing" vs. “sleeping"), high scores are assigned due to substantial syntactic overlap.
Motivated by these observations, we propose VC-Inspector, a novel factually grounded and reference-free model for inspecting the video caption quality. We leverage the recent advents of large multimodal models (LMMs) to handle long-context text reasoning and generalized video feature extraction, with the hypothesis that their established efficacy on visual-language joint reasoning can be applied to this task. While evaluation metrics are preferred to be lightweight, we employ the 3B/7B version of Qwen2.5-VL as a model-based evaluator and equip the model with the factual grounding ability through instruction tuning, where the model receives a video $V$ and a candidate caption $\hat{X}$, and generates an integer score ranging from 1 to 5. We also request the model to provide an explanation for the assigned score in the output, similar to [32], which makes the evaluator more interpretable and can serve as a supervision signal for factual grounding. The next section will cover how we systematically generate captions with incorrect factual elements, the corresponding quality estimates, and the explanations.
4 VC-Inspector: Instruction Tuning for Factually Grounded Evaluation
In video captioning, two primary factors that can degrade caption quality are: (1) the inclusion of objects not present in the video, and (2) the introduction of actions not depicted in the video. Our preliminary study indicates that existing metrics fall short in capturing these factual elements in the caption. To address this, we propose to empower LMMs with this ability through instruction tuning. However, it is nontrivial to acquire annotated samples for training. Supervised video caption datasets are already scarce due to the high cost of labeling, and these datasets generally only include correct captions, lacking those of lower quality. To obtain a considerable number of samples with diverse quality, we propose a systematic way to create a synthetic video caption dataset leveraging the internal knowledge of large language models (LLMs).
@@ 4 VC-Inspector: Instruction Tuning for Factually Grounded Evaluation
4.1 Data Generation
To create captions with incorrect elements, we employed Llama-3.3-70B-Instruct to alter the objects and actions in a ground truth caption from a supervised video caption dataset. Specifically, we created the ActivityNet-FG-It dataset for instruction tuning utilizing the ActivityNet [14] training set. The data generation pipeline is summarized in Figure 3.
@@ 4.1 Data Generation
After generating a pseudo caption, it is critical to assign it a reasonable quality score that is intuitive for humans so that it can better mirror humans’ expectations. To achieve this, one could utilize a visual-language model (VLM) to estimate this score; however, VLMs are not robust enough to produce a factually grounded score in a specified range without calibration, and the scoring quality is confined by their capabilities. Instead, we employ a deterministic scoring mechanism based on factual grounding, described as follows,
[…]
$=1-\frac{|{\cal R}|}{|{\cal O}|+|{\cal A}|}.$ (1)
Since $\cal{R}\subseteq{\cal O}\cup{\cal A}$ and $|{\cal O}\cup{\cal A}|=|{\cal O}|+|{\cal A}|$, the score will reside within the range of 0 and 1. It also guarantees that a caption with more incorrect objects/actions receives a lower score. We then convert the score to an integer in the 1-5 range for better interpretability during model training.
We repeated the aforementioned process to create 10 pseudo captions per ground truth caption, resulting in a total of 374K pseudo captions derived from 37,396 video-caption pairs. The prompts used to guide the generation process are detailed in the supplementary. The randomized replacement of objects and actions ensures coverage across the full range of the possible scores, as the number of replacements directly influences the semantic deviation from the original caption. However, it also naturally yields a skewed and non-uniform score distribution. To mitigate the potential bias during training, we applied a balanced sampling strategy, resulting in a refined subset of approximately 218K pseudo captions with uniform representation across five score categories. However, due to computational constraints, training with the full 218K instances would require multiple weeks of runtime. Therefore, we further sampled a 44K (8.8K for each label) caption subset from the balanced dataset for instruction tuning, which we refer to as ActivityNet-FG-It. This subset preserves category balance and maintains the diversity and structure of the original data, while offering a tractable size for experimentation ($\sim 32$ GPU hours using A100).
4.2 Training

### v2

# VC-Inspector: Advancing Reference-free Evaluation of Video Captions with Factual Analysis

@@ Abstract
We propose VC-Inspector, a lightweight, open-source large multimodal model (LMM) for reference-free evaluation of video captions, with a focus on factual accuracy. Unlike existing metrics that suffer from limited context handling, weak factuality assessment, or reliance on proprietary services, VC-Inspector offers a reproducible, fact-aware alternative that aligns closely with human judgments. To enable robust training and interpretable evaluation, we introduce a systematic approach for generating captions with controllable errors, paired with graded quality scores and explanatory annotations. Experiments show that VC-Inspector achieves state-of-the-art correlation with human judgments, generalizing across diverse domains (e.g., VATEX-Eval, Flickr8K-Expert, and Flickr8K-CF benchmarks) and revealing the potential for caption improvement.  [^1] https://dipta007.github.io/VC-Inspector/
VC-Inspector: Advancing Reference-free Evaluation
of Video Captions with Factual Analysis
Shubhashis Roy Dipta  1thanks: These authors contributed equally to this work.,  Tz-Ying Wu  2∗,  Subarna Tripathi  2 1University of Maryland, Baltimore County2Intel sroydip1@umbc.edu
@@ 1 Introduction
Video captions summarize salient objects and actions in videos, supporting downstream applications such as question answering, event localization, and retrieval P.J. and Kovoor (2024); Qian et al. (2024); Gabeur et al. (2020). Most evaluation metrics rely on reference-based protocols Papineni et al. (2002); Banerjee and Lavie (2005); Lin (2004); Vedantam et al. (2015); Zhang et al. (2020), which compare captions to human-written references. These approaches are costly to scale and often fail to capture semantic equivalence.
@@
Figure 1: Existing reference-free metrics like EMScore Shi et al. (2022) often fail to detect factual inaccuracies and lack a consistent scoring scale. VC-Inspector addresses these limitations by providing factually grounded, interpretable evaluations.
Evaluating captions for in-the-wild videos requires reference-free protocols that do not rely on ground truth captions, yet this remains underexplored. Existing metrics Shi et al. (2022); Sarto et al. (2023) typically measure visual-language alignment using pretrained multimodal embeddings Radford et al. (2021), but are limited by text encoder context length, and lack a consistent scoring scale, making interpretation difficult ( Fig.˜1). Recently, large proprietary models such as GPT-4o are used to rate captions on a fixed scale Tong et al. (2025), but these approaches depend heavily on prompt engineering and are not fully reproducible. Additionally, most prior methods are image-centric, rendering them suboptimal for video content.
In this work, we aim to develop a reference-free evaluation metric for video captions that removes dependency on human-annotated references while remaining robust and intuitive for humans. Our approach is grounded in factual accuracy, as factual elements like objects and actions are critical for video understanding. Ideally, a reliable metric should degrade scores proportionally to factual errors. For example, for a video showing a little girl sitting on a chair, the caption “A little girl is sleeping on a chair” should receive a higher score than “A woman is sleeping on a chair”, because the latter deviates more from the video content. However, existing metrics like EMScore Shi et al. (2022) often fail to capture even basic factual inaccuracies, such as incorrect objects/actions.
To address these limitations, we propose VC-Inspector, built on top of a lightweight, open-source large multimodal model (LMM) trained to assess caption quality based on factual correctness. Unlike previous metrics that provide only a score, VC-Inspector also generates explanations for its judgments, improving interpretability. One key challenge is the lack of captions with varying degrees of factual quality for training. We overcome this by introducing a novel data generation framework powered by a large language model (LLM), which controllably modifies factual elements in ground-truth captions from ActivityNet Krishna et al. (2017). This process yields ActivityNet-FG-It, a dataset of 44K examples for instruction tuning.  Fig.˜2 illustrates our data generation and training framework.
@@
We first evaluate VC-Inspector on two synthetic datasets, ActivityNet-FG-Eval and YouCook2-FG-Eval, generated using the same pipeline as our instruction-tuning dataset. Results show that VC-Inspector produces consistent quality estimates across datasets, showing versatility for diverse visual content and caption lengths. Next, we assess correlation with human judgments on VATEX-Eval Shi et al. (2022), a standard benchmark for video captioning metric evaluation. VC-Inspector outperforms prior reference-free metrics and even surpasses most reference-based metrics. Finally, we extend evaluation to image caption datasets, Flickr8K-Expert and Flickr8K-CF Hodosh et al. (2013), highlighting the strong generalization capability of our approach. Beyond the evaluation of the generated scores, we further show that the explanatory outputs of VC-Inspector not only enhance quality estimation but also contribute to caption refinement.
To summarize, this work makes the following contributions:
  • We introduce a scalable pipeline for synthesizing video captions with controllable factual errors, enabling large-scale training and evaluation, without costly human annotation.
  • We propose VC-Inspector, a fact-aware, reference-free video caption evaluator that jointly predicts quality scores and factual error explanations, enhancing interpretability and guiding caption improvement.
  • We demonstrate that VC-Inspector achieves high correlations with human judgments, outperforming existing reference-free methods and rivaling reference-based metrics across video and image caption benchmarks, showing strong cross-domain generalization.
@@ 2 Related Works
Traditional rule-based metrics such as BLEU Papineni et al. (2002), METEOR Banerjee and Lavie (2005), ROUGE Lin (2004), and CIDEr Vedantam et al. (2015) primarily capture syntactic similarity and fail to reflect semantic meaning. SPICE Anderson et al. (2016); Li et al. (2023) introduces semantic parsing of objects, attributes, and relations for both reference and candidate captions, but struggle with paraphrases. Embedding-based methods, such as BERTScore Zhang et al. (2020) and its variant Yi et al. (2020), which consider the intrinsic variance between multiple ground truth captions. More recently, CLAIR Chan et al. (2023a) explores an LLM-as-a-Judge approach for image caption evaluation, demonstrating stronger correlation with human judgments than the above metrics. However, all these metrics depend on reference captions for evaluation and ignore the visual content.
@@
To overcome the limitation of text-only metrics, image-aware approaches align captions with visual input. VIFIDEL Madhyastha et al. (2019) matches object names between images and captions, but is restricted to discrete categories and cannot capture motion dynamics in videos. Visual-language model (VLM) based metrics offer richer semantic alignment through continuous representations. ViLBERTScore Lee et al. (2020) extends BERTScore using ViLBERT Lu et al. (2019), but still requires comparison to reference captions. UMIC Lee et al. (2021) and CLIP-based metrics Hessel et al. (2022); Jiang et al. (2019); Cui et al. (2018); Sarto et al. (2023); Wada et al. (2024); Zeng et al. (2024); Lee et al. (2024); Shi et al. (2022); Liu and Wan (2023) relax this constraint by directly modeling image-caption semantic alignment via contrastive learning, though remain limited by the context length of their text encoders. Recently, Maeda et al. (2024) compares the candidate caption to the VLM-derived visual context with a LLM. However, these methods focus on static images and do not model the temporal dynamics, which limits their effectivenss for video caption evaluation.
In summary, text-based methods require costly, high-quality reference captions, while image-based extensions remain less effective for video content. These challenges underscore the need for reference-free video caption evaluation tailored to video inputs, a topic that is relatively less studied compared to text-based and image-caption evaluation. EMScore Shi et al. (2022) was an early attempt that supports evaluating video captions without a reference. Although it considers both frame-level and video-level embeddings, they are derived from an image-based encoder Radford et al. (2021), and its text encoder is constrained by short context length. PAC-S Sarto et al. (2023) and FactVC Liu and Wan (2023) augment EMScore with positive and negative data synthesis, respectively. While they share some similarity to this work, they consider only a single level of corruption, with binary (positive/negative) differentiation, whereas our method incorporates captions with varying degrees of quality, enabling a more nuanced evaluation. In addition, these metrics produce a single scalar score without offering explanations of their judgments, posing challenges for interpreting the quality assessment. More recently, G-VEval Tong et al. (2025) extends G-Eval Liu et al. (2023b) to video by stitching together three frames per video. However, it is unclear whether this image-based prompting generalizes to longer or more dynamic videos. Moreover, the dependence on large proprietary models such as GPT-4o limits its scalability and reproducibility for widespread use. In contrast, VC-Inspector finetunes a lightweight, open-source LMM to produce both evaluation scores and reasoning, offering a more interpretable and scalable solution.
3 Methodology
@@ 3.1 Overview
[…]
We argue that a reliable evaluation protocol should be factually grounded, interpretable, and scalable. It must reflect the correctness of factual elements (e.g., objects and actions) in the caption relative to the video. For example, captions with missing objects or incorrect actions should receive lower scores, and those with multiple errors should be penalized more severely. Furthermore, evaluation should be intuitive for humans and reproducible for broad adoption.
To address these limitations, we propose VC-Inspector, which leverages a lightweight, open-source large multimodal model (LMM) as the backbone for long-context reasoning and generalized video feature extraction. Our hypothesis is that the strong visual-language reasoning capability of LMMs can transfer effectively to this task. However, pretrained LMMs do not inherently produce factually grounded ratings and require targeted supervision (Table˜2). A key challenge is the scarcity of annotated captions exhibiting diverse degrees of factual accuracy for effective training. To overcome this, we introduce a novel data generation pipeline powered by large language models (LLMs) that systematically synthesizes candidate captions along with quality estimates and factual error explanations. The overall data generation and training framework is illustrated in Fig.˜2.
3.2 Data Generation
To create captions with controlled factual errors, we employ Llama-3.3-70B-Instruct to systematically alter objects and actions in ground truth captions from a supervised video caption dataset. Using this approach, we construct ActivityNet-FG-It, an instruction-tuning dataset derived from the ActivityNet Krishna et al. (2017) training set. The data generation pipeline is illustrated in Fig.˜3.
Caption generation.
Given a ground truth caption $X$, we first prompt the LLM to extract the set of objects ${\cal O}=\{o_{1},...,o_{M}\}$ and actions ${\cal A}=\{a_{1},...,a_{N}\}$. We then randomly sample $K\sim Unif(0,M)$ objects and $L\sim Unif(0,N)$ actions to replace, forming a subset $\cal{R}\subseteq{\cal O}\cup{\cal A}$, where $|{\cal R}|=K+L$ and $Unif(a,b)$ denotes a discrete uniform distribution over integers $\{a,...,b\}$.
[…]
This process also yields a factual error explanation for $\tilde{X}$, enabling fine-grained supervision. The detailed prompts and explanation formats are provided in App.˜D.
@@ 3.2 Data Generation
Scoring.
After generating a pseudo caption, assigning an intuitive quality score is essential to align with human expectations. While pretrained embeddings like visual-language model (VLM) can estimate this score, they lack a fixed scale, are constrained by context length, and do not guarantee factual grounding. Instead, we adopt a deterministic scoring mechanism based on factual accuracy:
@@
$=1-\frac{|{\cal R}|}{|{\cal O}|+|{\cal A}|}.$
Since $\cal{R}\subseteq{\cal O}\cup{\cal A}$ and $|{\cal O}\cup{\cal A}|=|{\cal O}|+|{\cal A}|$, the score lies on a fixed scale between 0 and 1, ensuring captions with more incorrect elements receive lower scores. Since LLMs are inherently unreliable when comparing floating-point values (Spathis and Kawsar, 2023), we discretize the score into a 1-5 range ($score=round(score\times 4+1)$), consistent with standard human annotation protocols (e.g., Likert-style ratings in VATEX-Eval). While this may induce minor loss of information (e.g., $0.87\rightarrow 4$ and $0.88\rightarrow 5$), this conversion preserves ordinal relationships and error severity while maintaining the underlying deterministic structure.
Moreover, VC-Inspector is jointly trained to produce textual explanations that explicitly identify incorrect objects and actions, providing fine-grained supervision beyond the scalar score.
Post-processing.
We repeat the caption generation and scoring process to create 10 pseudo captions per ground truth caption, yielding 374K pseudo captions derived from 37,396 video-caption pairs. The randomized replacement of objects and actions ensures coverage across the full range of the possible scores, as the number of replacements directly influences the semantic deviation. However, this naturally results in a skewed and non-uniform score distribution. To mitigate the potential bias during training, we apply a balanced sampling strategy, resulting in a refined subset of approximately 218K pseudo captions with uniform representation across five score categories.
Due to computational constraints, training on the full 218K instances would require multiple weeks. Therefore, we further sample 44K captions (8.8K per label) from the balanced dataset for instruction tuning, which we refer to as ActivityNet-FG-It. This subset preserves category balance and maintains the diversity, while offering a tractable size for experimentation ($\sim 32$ GPU hours using A100). Importantly, although we use Llama as the generator and perform subsampling as a practical design choice, our data generation pipeline is model-agnostic and applies across model scales and caption datasets, enabling scalable dataset construction at arbitrary sizes.
3.3 Training of VC-Inspector